# Non-thermal emission from mildly relativistic tidal ejecta of compact objects merger

### v1

@@ 3 Semi-analytic calculation
A semi-analytic approach for calculating the expected synchrotron emission from a toroidal ejecta with a velocity profile given by Eq. (1) is presented below. This approach is valid for as long as the reverse shock propagates through the ejecta steep, fast tail. After the reverse shock crosses the fast tail, a full 2D calculation is needed for a reliable calculation of the synchrotron emission. The semi-analytic analysis code will be shared following a reasonable request to the corresponding author, offering a computation method for the approximated dynamics, anticipated flux, and the radio image on the sky.
@@ 3.1 Dynamics
The mass of a toroid ejecta, $M$, with an opening angle of $\theta_{\text{open}}$ from the equatorial plane corresponds to $\sin{\theta_{\text{open}}}M_{\text{iso}}$, where $M_{\text{iso}}$ is the isotropic equivalent mass. The ejecta propagates through uniform ISM density, forming a forward-reverse shock structure; see Fig. 1. As long as the reverse shock did not cross the fast tail (the steep, fast part of the ejecta), the total internal energy (and thus, the emissivity) of the system is dominated by the freshly shocked material within the initial opening angle because of the following argument: due to the steep decline of the ejecta profile ($s_{\text{ft}}\geq 5$) the injected energy (and momentum) to the shocked medium increases faster with decreasing velocity (as the ejecta mass increases faster with decreasing velocity). The swept ISM mass is dominated by the freshly shocked material since the shock is propagating radially and, due to momentum conservation, the momentum of the freshly shocked material is dictated by added momentum (to the shocked medium) from the ejecta at the position of the reverse shock, which is in the radial direction. We conclude the shocked plasma lateral expansion should have a limited effect on the observed emission before the reverse shock crosses the steep, fast tail. Thus, we approximate the ejecta deceleration dynamics using the dynamics of stratified spherical ejecta with a corresponding mass of $M_{\text{iso}}=M/(\sin{\theta_{\text{open}}})$ ).
@@ 3.3 Intensity and flux
Using the approximated dynamics mentioned above, we obtain the flow profiles of $e(R,t),\rho^{\prime}(R,t)$ and $\gamma(R,t)$ (internal energy density, proper mass density, and Lorentz factor correspondingly as a function of radius and time in the rest frame of the ISM) behind the shocks front, with different values for the forward and reverse shocks. We approximate the shocked plasma as two separate uniform layers behind the shocks with the flow profiles described above. The thickness of these layers is determined by conservation of mass both for forward and reverse shocks. We employ a phenomenological approach to describe synchrotron emissivity: the fraction of energy carried by electrons is denoted by $\varepsilon_{e}$, and the fraction of energy carried by magnetic fields is denoted by $\varepsilon_{B}$. We assume a power-law electron distribution within the shocked layers, $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$. Following the assumption that the emission is dominated by shocked plasma within the initial opening angle, we restrict the emissivity to $\theta_{\text{open}}$ of the ejecta. Mathematically, this is expressed as:
@@
where $j^{\prime}_{\nu}\left(e,\rho,\gamma\right)$ is the emissivity as a function of the flow profiles, and the prime () notation indicates proper frame quantities. Subsequently, the emissivity and intensity are computed following the methodology outlined in Sadeh et al. (2024), where we only considered frequencies above the self-absorption frequency and below the cooling frequency, $\nu_{a}<\nu<\nu_{c}$. To calculate the flux, we consider the fact that the azimuthal observer symmetry breaks in the case of $\theta_{\text{v}}\neq 0$, so we calculate each azimuthal angle separately by
@@
For the light curve temporal cutoff, we use the peak time estimation from Sadeh et al. (2023),
@@
and $F^{\text{iso}}_{\nu,\rm peak}$ is the peak flux in the equivalent isotropic case.
@@ 4.1 Numerical set-up
which is in very good agreement with the EoS of mildly relativistic fluid provided by Synge (1957). The $1.3$ factor is added for better agreement with the exact solution in the range of $20<e/\rho^{\prime}c^{2}<$ 2 (corresponding to $31<\gamma<$ 3); see appendix B. Our simulations were performed in a 2D cylindrical coordinate system $(r,z)$ assuming axial symmetry, and we adopted a Courant–Friedrichs–Lewy (CFL) number of $0.5$. The initial grid prioritizes resolving the ejecta by placing a higher concentration of cells within it, with a coarser spacing outside. Typically, the simulation begins with $500\times 500$ cells, focusing most cells within the ejecta. AMR dynamically adjusts the resolution throughout the simulation, optimizing cell distribution along both the $r$ and $z$ axes to effectively capture regions of significant variation in pressure, density, or Lorentz factor. The initial pressure in all of the cells is set to $P=10^{-10}\rho_{\text{ISM}}c^{2}$ (where $\rho_{\text{ISM}}$ is the ISM mass density), and the computational domain boundaries are set at $r=10^{18}$cm and $z=10^{18}$cm. We systematically refined the initial spatial grid and temporal steps to assess convergence. We also iteratively increased the number of times the regridding scheme is allowed to multiply the number of cells within a time step. This process continues until no significant changes are observed in the hydrodynamics and emergent flux. Reflective boundary conditions were implemented at both the symmetry axis ($r=0$) and the symmetry plane ($z=0$) to account for symmetry considerations. Our numerical calculations adopt initial conditions comprising toroid ejecta, as illustrated in Fig. 1. The ejecta is characterized by a flow profile described in Eq. (1). The initial, $t=t_{0}$, radius of a mass shell $M(>\gamma\beta)$ is determined by $r_{0}=\beta ct_{0}$. The expanding ejecta is initially (at $t=t_{0}$) embedded (at $r>ct_{0}$) in a static, uniform, cold (negligible pressure) gas.
@@ 4.2 Numerical light curve
From the 2D grid obtained via RELDAFNA, we construct a 3D grid by subdividing cells into segments that represent different azimuthal angles such that the number of segments is appropriate for achieving flux convergence. By post-processing the hydrodynamic profiles, we calculate specific emissivity, $j^{\prime}_{\nu}(\vec{r},t)$ (following Rybicki & Lightman, 1979; Sadeh et al., 2024, appendix B, for frequency below the cooling frequency and above the synchrotron and self-absorption frequencies, $\nu_{a},\nu_{m}<\nu<\nu_{c}$), in the plasma frame for each cell within the 3D grid at a discrete set of lab times. Each fluid element is assumed to exist for a period of $\Delta t$ interval between time steps. This calculation adopts a phenomenological approach for the magnetic field and electron energy distribution, assuming fractions $\varepsilon_{B}$ and $\varepsilon_{e}$ of internal energy density for magnetic fields and electrons, respectively. Furthermore, we assume a power-law energy distribution for electrons, $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$. The flux contribution from each cell for every observed time is computed as follows,
$\Delta F_{\nu}\left(t_{\text{obs}}=t-\frac{R\cos{\xi}}{c}\right)=\frac{j^{\prime}_{\nu}(t)}{\gamma(t)^{2}(1-\beta(t)\cos{\xi_{v}})^{2}}\cdot\frac{\Delta V}{D^{2}},$ (10)
where $R$ is the distance from the origin, $\Delta V$ is the cell’s volume, $D$ is the luminosity distance, $t$ is the time in the lab frame, $t_{\text{obs}}$ is the observer time, $\xi$ is the angle between the cell and the line of sight and $\xi_{v}$ is the angle between the cell velocity direction and the line of sight. $\xi$ and $\xi_{v}$ pose different values for different observation angles following Eq. (2). The total observed flux per unit frequency and observed time $t_{\text{obs}}$ is obtained by summing the different contributions $\Delta F_{\nu}$ from all of the fluid elements at all the discretized times.
@@ 5 Validation of the semi-analytic method
In 5.1, we verify our estimation regarding the dominant contribution to the internal energy by the freshly shocked material within the initial opening angle of the ejecta. This verification holds true as long as the reverse shock hasn’t yet crossed the ejecta fast tail. In 5.2, we compare the light curves produced by the full 2D numerical calculation to our semi-analytic scheme to verify its validity. Finally, in 5.3, we compare the flux angular distribution obtained by our semi-analytic method to the full 1D numerical calculation of a spherical ejecta case to approve validation.
@@ 5.1 Hydrodynamics
In our semi-analytic method, we assume that the internal energy is dominated by material within the initial opening angle of the ejecta. We ran numeric calculations with various initial conditions to test our assumption. The opening angle as a function of lab time is defined as the angle from the equatorial plane within which $90\%$ of the internal energy, excluding rest mass energy, is included. A few examples are given in Fig. 2. We indeed find that the contribution to the internal energy is dominated by the material within the initial opening angle (up to $\sim 30\%$ corrections) as long as the reverse shock doesn’t cross the ejecta fast tail.
Figure 2: The opening angle (relative to the ejecta initial opening angle) within which $90\%$ of the internal energy, excluding rest mass energy, is included for various toroid ejecta parameters; $\{s_{\text{ft}},s_{\text{KN}},\gamma_{0},n,M_{R},\theta_{\text{open}}\}=\{6,1.5,1.2,3\times 10^{-2}\text{cm}^{-3},10^{-8}M_{\odot},30^{\circ}\}$ /$/\{7,1.5,1.15,3\text{cm}^{-3},10^{-6}M_{\odot},20^{\circ}\}$ /$/\{8,1.5,1.1,3\text{cm}^{-3},10^{-6}M_{\odot},15^{\circ}\}$ /$/\{9,1.5,1.15,3\times 10^{-1}\text{cm}^{-3},10^{-7}M_{\odot},25^{\circ}\}$.
@@ 5.2 Light curves
We consider only frequencies that are between the self-absorption frequency/peak frequency to the cooling frequency, $\nu_{a},\nu_{m}<\nu<\nu_{c}$, since the radio observations are expected to be in this part. In Fig. 3, we compare the light curve obtained semi-analytically (for as long as the reverse shock doesn’t cross the fast tail, § 3) with the full 2D numerical calculation (§ 4). The agreement is within $10$’s of percent for a wide range of relevant values of $\{\beta_{0},s_{\rm ft},\theta_{\text{v}},\theta_{\text{open}}\}$. As expected, a $\sim 20\%$ delay is observed for $\theta_{\text{v}}=0^{\circ}$.
Figure 3: A comparison between the flux calculated semi-analytically in (solid lines) and the flux calculated in the full numerical calculation (dashed lines), for various parameters $\{s_{\text{ft}},s_{\text{KN}},\gamma_{0},n,M_{R},\varepsilon_{e},\varepsilon_{B},p,\theta_{\text{open}}\}=\{6,1.5,1.2,3\times 10^{-2}\text{cm}^{-3},10^{-8}M_{\odot},10^{-1},10^{-1},2.1,30^{\circ}\}$ /$/\{7,1.5,1.15,3\text{cm}^{-3},10^{-6}M_{\odot},10^{-1},10^{-1},2.2,20^{\circ}\}$ /$/\{8,1.5,1.1,3\text{cm}^{-3},10^{-6}M_{\odot},10^{-1},10^{-1},2.3,15^{\circ}\}$ /$/\{9,1.5,1.15,3\times 10^{-1}\text{cm}^{-3},10^{-7}M_{\odot},10^{-1},10^{-1},2.4,25^{\circ}\}$. We used a distance of $100$Mpc and a frequency of $1$GHz.

### v2

@@ 3 Semi-analytic calculation
A semi-analytic approach for calculating the expected synchrotron emission from a toroidal ejecta with a velocity profile given by Eq. (1) is presented below. This approach is valid as long as the reverse shock propagates through the ejecta steep, fast tail. After the reverse shock crosses the fast tail, a full 2D calculation is needed for a reliable calculation of the synchrotron emission. The semi-analytic analysis code offers a computation method for the approximated dynamics, anticipated flux, and the radio image on the sky.
@@ 3.1 Dynamics
The toroidal ejecta has mass, $M=M_{\text{iso}}\sin{\theta_{\text{open}}}$ ), where $M_{\text{iso}}$ is the isotropic equivalent mass and $\theta_{\text{open}}$ is the ejecta opening angle. The ejecta propagates through uniform ISM density, forming a forward-reverse shock structure, as shown in Fig. 1. As long as the reverse shock did not cross the steep fast tail, the total internal energy (and thus, the emissivity) of the system is dominated by the freshly shocked material within the initial opening angle because of the following argument: Given the steep ejecta profile with a power-law index of $s_{\text{ft}}\geq 5$, the energy and momentum injected into the shocked medium increase rapidly with decreasing velocity. This is because the mass of the ejecta, which scales as $M(>\gamma\beta)\propto(\gamma\beta)^{-s_{\text{ft}}}$, grows more quickly as velocity decreases. Consequently, the freshly shocked material, which is continuously energized by this injection, dominates the internal energy of the system. As the shock front propagates radially through the interstellar medium (ISM), the swept-up ISM mass is primarily composed of this freshly shocked material. The unshocked ejecta, confined within an opening angle $\theta_{\text{open}}$, also moves radially, injecting additional energy and momentum into the shocked plasma. Due to conservation of momentum, the momentum of the freshly shocked material is dictated by this added momentum from the ejecta at the position of the reverse shock. We conclude the shocked plasma lateral expansion should have a limited effect on the observed emission before the reverse shock crosses the steep, fast tail. For larger values of $s_{\text{ft}}$, this effect becomes more pronounced, resulting in a less substantial lateral expansion. Thus, we approximate the ejecta deceleration dynamics using the dynamics of stratified spherical ejecta with a corresponding mass of $M_{\text{iso}}=M/(\sin{\theta_{\text{open}}})$ ).
@@ 3.3 Intensity and flux
Using the approximated dynamics mentioned above, we obtain the flow profiles behind the shocks front: the internal energy density $e(R,t)$, the proper mass density $\rho^{\prime}(R,t)$, and the Lorentz factor $\gamma(R,t)$, all as a function of the shock radius and time in the rest frame of the ISM. We approximate the shocked plasma as two separate uniform layers behind the shocks with different flow profiles for the two different layers. The thickness of these layers is determined by conservation of mass both for forward and reverse shocks. We employ a phenomenological approach to describe synchrotron emissivity: the fraction of energy carried by electrons is denoted by $\varepsilon_{e}$, and the fraction of energy carried by magnetic fields is denoted by $\varepsilon_{B}$. We assume a power-law electron distribution within the shocked layers, $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$. Following the assumption that the emission is dominated by shocked plasma within the initial opening angle, we restrict the emissivity to within $\theta_{\text{open}}$. Mathematically, this is expressed as:
@@
where $j^{\prime}_{\nu}\left(e,\rho,\gamma\right)$ is the emissivity as a function of the flow profiles, and the prime () notation indicates proper frame quantities. Subsequently, the emissivity and intensity are computed by the methodology outlined in Sadeh & Waxman (2024): $j_{\nu}$ is defined by following Rybicki & Lightman (1979), and $I_{\nu}$ is calculated by integrating over the $j_{\nu}$ while taking into account arrival time effects and relativistic effects. We only consider frequencies above the self-absorption frequency and below the cooling frequency, $\nu_{a}<\nu<\nu_{c}$, since the bulk of the radio to X-ray observations are expected to be in this part of the spectrum (Sadeh & Waxman, 2024). To calculate the flux, we consider the fact that the azimuthal observer symmetry breaks in the case of $\theta_{\text{v}}\neq 0$, so $I_{\nu}$ varies for different $\varphi$ values. We integrate over the contribution from each azimuthal angle by
@@
where $R$ is the shock radius. For the light curve temporal cutoff, we use the peak time estimation from Sadeh et al. (2023),
@@
$l_{p}=\frac{p-2}{1-(\gamma_{\text{max}}/\gamma_{\text{min}})^{2-p}},\quad l_{p}\xrightarrow[p\to 2]{}\frac{1}{\ln(\gamma_{\text{max}}/\gamma_{\text{min}})},$ (10)
and $F^{\text{iso}}_{\nu,\rm peak}$ is the peak flux in the equivalent isotropic case. $f_{\rm ft}(p)$ deviates from unity by less than $30\%$ for $2.52<p<$ 2.5 (Sadeh et al., 2023).
@@ 4.1 Numerical set-up
which is in very good agreement with the EoS of mildly relativistic fluid provided by Synge (1957). The $1.3$ factor is added for better agreement with the exact solution in the range of $20<e/\rho^{\prime}c^{2}<$ 2 (corresponding to $31<\gamma<$ 3); see Appendix B. Our simulations were performed in a 2D cylindrical coordinate system $(r,z)$ assuming axial symmetry, and we adopted a Courant–Friedrichs–Lewy (CFL) number of $0.5$, which provided efficient convergence (physical justified as long as it is $\leq 1$). The initial grid prioritizes resolving the ejecta by placing a higher concentration of cells within it, with a coarser spacing outside. Typically, the simulation begins with $500\times 500$ cells, focusing most cells within the ejecta. AMR dynamically adjusts the resolution throughout the simulation, optimizing cell distribution along both the $r$ and $z$ axes to effectively capture regions of significant variation in pressure, density, or Lorentz factor. The initial pressure in all of the cells is set to $P=10^{-10}\rho_{\text{ISM}}c^{2}$ (where $\rho_{\text{ISM}}$ is the ISM mass density), and the computational domain boundaries are set at $r=10^{18}$cm and $z=10^{18}$cm. We systematically refined the initial spatial grid and temporal steps to assess convergence. We also iteratively increased the number of times the regridding scheme is allowed to multiply the number of cells within a time step. This process continues until no significant changes are observed in the hydrodynamics and emergent flux. Reflective boundary conditions were implemented at both the symmetry axis ($r=0$) and the symmetry plane ($z=0$) to account for symmetry considerations. Our numerical calculations adopt initial conditions comprising toroid ejecta, as illustrated in Fig. 1. The ejecta is characterized by a flow profile described in Eq. (1). The initial radius (at $t=t_{0}$) of a mass shell $M(>\gamma\beta)$ is determined by $r_{0}=\beta ct_{0}$. The expanding ejecta is initially embedded at $r>ct_{0}$ in a static, uniform, cold gas with negligible pressure. See Fig. 2 for an example of a density map snapshot calculated with RELDAFNA.
Figure 2: An example of a 2D density profile of a toroidal outflow propagating into a uniform interstellar medium calculated with RELDAFNA (the azimuthal symmetry axis is located at $x=0$). The initial time of the calculation is $10^{5}$s, and the snapshot is taken at $t=1.35\times 10^{5}$s. The density scale is logarithmic. The initial conditions of the outflow are $\{s_{\text{ft}},s_{\text{KN}},\beta_{0},n,M_{R},\theta_{\text{open}}\}=\{6,1.5,0.3,3\times 10^{2}\text{cm}^{-3},10^{-4}M_{\odot},20^{\circ}\}$.
@@ 4.2 Numerical light curve
From the 2D grid obtained via RELDAFNA, we construct a 3D grid by subdividing cells into segments that represent different azimuthal angles such that the number of segments is appropriate for achieving flux convergence. By post-processing the hydrodynamic profiles, we calculate specific emissivity, $j^{\prime}_{\nu}(\vec{r},t)$ (following Rybicki & Lightman, 1979; Sadeh & Waxman, 2024, Appendix B), for frequency below the cooling frequency and above the synchrotron and self-absorption frequencies, $\nu_{a},\nu_{m}<\nu<\nu_{c}$. The emissivity is calculated in the plasma proper frame for each cell within the 3D grid at a discrete set of lab times. Each fluid element is assumed to exist for a period of $\Delta t$ interval between time steps. This calculation adopts a phenomenological approach for the magnetic field and electron energy distribution, assuming fractions $\varepsilon_{B}$ and $\varepsilon_{e}$ of internal energy density for magnetic fields and electrons, respectively. Furthermore, we assume a power-law energy distribution for electrons, $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$. The flux contribution from each cell for every observed time is computed as follows,
$\Delta F_{\nu}\left(t_{\text{obs}}=t-\frac{R\cos{\xi}}{c}\right)=\frac{j^{\prime}_{\nu}(t)}{\gamma(t)^{2}(1-\beta(t)\cos{\xi_{\beta}})^{2}}\cdot\frac{\Delta V}{D^{2}},$ (12)
where $R$ is the distance from the origin, $\Delta V$ is the cell’s volume, $D$ is the luminosity distance, $t$ is the time in the lab frame, $t_{\text{obs}}$ is the observer time, $\xi$ is the angle between the cell and the line of sight and $\xi_{\beta}$ is the angle between the cell velocity direction and the line of sight. $\xi$ pose different values for different observation angles following Eq. (2). Similarly, $\xi_{\beta}$ shares the same property by considering the following transformation, $\xi\rightarrow\xi_{\beta},\theta\rightarrow\theta_{\beta}$ (where $\theta_{\beta}$ is the angle between the velocity direction and the symmetry axis), to Eq. (2). The total observed flux per unit frequency and observed time $t_{\text{obs}}$ is obtained by summing the different contributions $\Delta F_{\nu}$ from all of the fluid elements at all the discretized times.
@@ 5 Validation of the semi-analytic method
In § 5.1, we verify our estimation regarding the dominant contribution of the freshly shocked material within the initial opening angle of the ejecta to the internal energy. This verification holds true as long as the reverse shock hasn’t yet crossed the ejecta fast tail. In § 5.2, we compare the light curves produced by the full 2D numerical calculation with our semi-analytic scheme to verify its validity. Finally, in § 5.3, we compare the flux angular distribution obtained by our semi-analytic method to the full 1D numerical calculation of a spherical ejecta case to test its validity.
@@ 5.1 Hydrodynamics
In our semi-analytic method, we assume that the internal energy is dominated by material within the initial opening angle of the ejecta. We ran numeric calculations with various initial conditions to test our assumption. The opening angle as a function of lab time is defined as the angle from the equatorial plane within which $90\%$ of the internal energy, excluding rest mass energy, is included. A few examples are given in Fig. 3. We consider various parameter options since all of the following parameters $\{s_{\text{ft}},n,M_{R},\}$ affect the ejecta dynamics, where $M_{R}\equiv M(\gamma\beta>1)=M_{0}(\gamma_{0}\beta_{0})^{s_{\rm ft}}$, is the "relativistic mass" with momentum $\gamma\beta>1$. We consider only typical values for the opening angle, $\theta_{\text{open}}=20^{\circ}$, and the velocity cutoff, $\beta_{0}=0.3-0.4$, that are consistent with the ejected material in merger simulations. We indeed find that the contribution to the internal energy is dominated by the material within the initial opening angle (up to $\sim 30\%$ corrections) as long as the reverse shock doesn’t cross the ejecta fast tail. An evident trend in Fig. 3 is that larger values of $s_{\text{ft}}$ lead to a less pronounced lateral expansion. This is because, at larger $s_{\text{ft}}$, the freshly shocked material makes up a larger fraction of the shocked mass, as explained in § 3.1.
Figure 3: The opening angle (relative to the ejecta initial opening angle) within which $90\%$ of the internal energy, excluding rest mass energy, is included. We consider few sets of toroidal ejecta parameters that affect the outflow dynamics: $\{s_{\text{ft}},\beta_{0},\theta_{\text{open}}\}=\{6,0.3,20^{\circ}\}$ /$/\{7,0.35,20^{\circ}\}$ /$/\{8,0.4,20^{\circ}\}$, and $\{n,M_{R}\}=A:\{3\times 10^{2}\text{cm}^{-3},10^{-4}M_{\odot}\},B:\{3\text{cm}^{-3},10^{-6}M_{\odot}\}$. The results are plotted for as long as the reverse shock propagates through the ejecta fast tail.
@@ 5.2 Light curves
We consider only frequencies that are between the self-absorption frequency/peak frequency to the cooling frequency, $\nu_{a},\nu_{m}<\nu<\nu_{c}$, since the radio observations are expected to be in this part. In Fig. 4, we compare the light curve obtained semi-analytically (for as long as the reverse shock doesn’t cross the fast tail, § 3) with the full 2D numerical calculation (§ 4). The agreement is within $10$’s of percent for a range of relevant values of $\{n,M_{R},\beta_{0},s_{\rm ft},\theta_{\text{v}},\theta_{\text{open}},p\}$, along with $\varepsilon_{e}=10^{-1}$ and $\varepsilon_{B}=10^{-2}$, which modify the light curve based on a known analytical dependence, and $s_{\text{KN}}=1.5$, which alters the light curve after the peak. As expected, a $\sim 20\%$ delay is observed for $\theta_{\text{v}}=0^{\circ}$.
Figure 4: A comparison between the flux calculated semi-analytically in (solid lines) and the flux calculated in the full numerical calculation (dashed lines), for various parameters. Left panel: $\{s_{\text{ft}},\beta_{0},p,\theta_{\text{v}}\}=\{7,0.5,2.2,0^{\circ}/30^{\circ}/60^{\circ}\}$ and $\{n,M_{R}\}=\{3\text{cm}^{-3},10^{-6}M_{\odot}\}$. Right panel: $\{s_{\text{ft}},\beta_{0},p,\theta_{\text{v}}\}=\{6,0.3,2.1,30^{\circ}/90^{\circ}\}$ /$/\{7,0.35,2.2,30^{\circ}/90^{\circ}\}$ and $\{n,M_{R}\}=\{3\times 10^{2}\text{cm}^{-3},10^{-4}M_{\odot}\}$. We also used $\{\varepsilon_{e},\varepsilon_{B},\theta_{\text{open}},s_{\text{KN}}\}=\{10^{-1},10^{-2},20^{\circ},1.5\}$, a distance of $100$Mpc, and a frequency of $1$GHz.